OpenStudy (erinweeks):

What are all the real zeros of y = (x - 12)3 - 10?

OpenStudy (anonymous):

Try to expand the function then factor.

OpenStudy (erinweeks):

A. \[x = \sqrt[3]{10 + 12}\] B. \[x = \sqrt[3]{10} + 12\] C. x = \[\sqrt[3]{-10} + 12\] D. \[x = \sqrt[3]{12} -10\]

OpenStudy (erinweeks):

how do i expand it could you help?

OpenStudy (anonymous):

Wait, nevr mind about that. You can try and test each choice. One of them will cause the entire expression to be zero.

OpenStudy (erinweeks):

how do i test them?

OpenStudy (erinweeks):

@jim_thompson5910 please help.

jimthompson5910 (jim_thompson5910):

\[\Large y = (x - 12)^3 - 10\] \[\Large 0 = (x - 12)^3 - 10\] \[\Large 0+10 = (x - 12)^3 - 10+10\] \[\Large 10 = (x - 12)^3\] \[\Large (x - 12)^3 = 10 \] See what to do from here?

OpenStudy (erinweeks):

yes. the answer has to equal zero or what ... like what numbers do i plug in

jimthompson5910 (jim_thompson5910):

your next step is to take the cube root of both sides

jimthompson5910 (jim_thompson5910):

\[\Large y = (x - 12)^3 - 10\] \[\Large 0 = (x - 12)^3 - 10\] \[\Large 0+10 = (x - 12)^3 - 10+10\] \[\Large 10 = (x - 12)^3\] \[\Large (x - 12)^3 = 10 \] \[\Large \sqrt[3]{(x - 12)^3} = \sqrt[3]{10} \] \[\Large x - 12 = \sqrt[3]{10} \] then what?

OpenStudy (erinweeks):

subtract 12 ? im not sure.

jimthompson5910 (jim_thompson5910):

actually add 12 to both sides

jimthompson5910 (jim_thompson5910):

\[\Large y = (x - 12)^3 - 10\] \[\Large 0 = (x - 12)^3 - 10\] \[\Large 0+10 = (x - 12)^3 - 10+10\] \[\Large 10 = (x - 12)^3\] \[\Large (x - 12)^3 = 10 \] \[\Large \sqrt[3]{(x - 12)^3} = \sqrt[3]{10} \] \[\Large x - 12 = \sqrt[3]{10} \] \[\Large x - 12+12 = \sqrt[3]{10}+12 \] \[\Large x = \sqrt[3]{10}+12 \]

jimthompson5910 (jim_thompson5910):

so to sum things up: you plug in y = 0, then you solve for x

jimthompson5910 (jim_thompson5910):

to get \[\Large x = \sqrt[3]{10}+12 \]

OpenStudy (erinweeks):

okay i get it a little better. thank you Jim

jimthompson5910 (jim_thompson5910):

np